suppose that a company manufactures flash drives has a revenue of given by: R(x)10x-.001x^2
where the production output in 1 week is x flash drives. If production is increasing at a rate of 650 flash drives per week when production is 6000 flash drives find the rate of increase in revenue.
1
Expert's answer
2019-12-16T09:12:50-0500

In this we have to find out the rate of increase in revenue which will be calculated by finding out the derivative of the R(x)

so we have been given R(x)=10x-0.001x2

where the production output in 1 week is x flash drives

now, differentiating R(x) with respect to t

we get , R'(x)=10*x'-0.001*2*x*x' ......(1)

where x' is the the derivative of x with respect of t which is basically the rate of increase of production per week

given in the question x= 6000 and x'= 650

on substituting the values of x and x' in equation (1)

we get R'(x)=10*650-0.001*2*6000*650

R'(x)=6500-12*650

R'(x)=-1300

hence the value of R'(x) comes out to be negative which means rate of revenue will decrease

rate of increase in revenue = -1300